OpenStudy (anonymous):

Rewrite the expression (cos^2x)/(1+sinx) so that there is no fraction.

OpenStudy (saifoo.khan):

\[\frac{1-\sin^2x}{1+\sin^2x}\] Remember difference of squares? \[\frac{(1-\sin x)(1+\sin x)}{1+\sin x}\] can you solve now?

OpenStudy (anonymous):

is it 1/sinx = cscx?

OpenStudy (saifoo.khan):

1+sin x cancels out from both numerator and denominator and so leaving you behind with 1-sin x

OpenStudy (anonymous):

Got it! Thanks!